# Second- and third-order properties of multidimensional Langevin equations

### v9

@@ 10. Appendix A
(374) $L(w,w^{*})=-L(w^{*},w)=-(L(w,w^{*}))^{*},$
which is purely imaginary, but not necessarily non-zero. If $x$ is real, we also have that $\widetilde{L}(x,w,w^{*})$ is purely imaginary. For judging quantitative significance, we need to treat $w$ and $w^{*}$ as separate variables. Otherwise, we use the same procedure. For example, suppose we have a complex variable $w$ obeying the symmetry $w\to e^{i\phi}w$. We then have, for the “ensemble variance”:
@@
(377) $=\left({\Delta_{\textrm{exp\textendash theo}}\langle x^{2}\rangle}\right)^{2}+\left({\Delta_{\textrm{exp\textendash theo}}\langle y^{2}\rangle}\right)^{2}+2\left({\Delta_{\textrm{exp\textendash theo}}\langle x^{2}\rangle}\right)\left({\Delta_{\textrm{exp\textendash theo}}\langle y^{2}\rangle}\right)$
(378) $\sim\langle x^{2}\rangle^{2}+\langle y^{2}\rangle^{2}+2\langle xy\rangle^{2}$
(379) $=\frac{\langle ww^{*}\rangle^{2}}{2},$
@@
For example, if we have complex variables $w_{1},w_{2}$ obeying the symmetry $w_{i}\to e^{i\phi}w_{i}$ (simultaneous transformation of $w_{1},w_{2}$), then we have:
@@ 13. Acknowledgments
Y.I.L. acknowledges support from the McGill University Department of Physics.

### v11

@@ 11. Appendix A
(406) $L(w,w^{*})=-L(w^{*},w)=-(L(w,w^{*}))^{*}$
($w^{*}$ being the complex conjugate of $w$), which is purely imaginary, but not necessarily zero. If $x$ is real, we also have that $\widetilde{L}(x,w,w^{*})$ is purely imaginary. For judging quantitative significance, we need to treat $w$ and $w^{*}$ as separate variables. Otherwise, we use the same procedure. For example, suppose we have a complex variable $w$ obeying the symmetry $w\to e^{i\phi}w$ for any real $\phi$. We then have, for the “ensemble variance”:
@@
$=\left({\Delta_{\textrm{exp\textendash theo}}\langle x^{2}\rangle}\right)^{2}+\left({\Delta_{\textrm{exp\textendash theo}}\langle y^{2}\rangle}\right)^{2}+2\left({\Delta_{\textrm{exp\textendash theo}}\langle x^{2}\rangle}\right)\left({\Delta_{\textrm{exp\textendash theo}}\langle y^{2}\rangle}\right)$
$\sim\langle x^{2}\rangle^{2}+\langle y^{2}\rangle^{2}+2\langle xy\rangle^{2}$
$=\frac{\langle ww^{*}\rangle^{2}}{2},$
@@
For example, if we have complex variables $w_{1},w_{2}$ obeying the symmetry $w_{i}\to e^{i\phi}w_{i}$ for any real $\phi$ (where $w_{1},w_{2}$ are simultaneously transformed), then we have:
@@
Also, for complex Gaussian variables $z_{1},z_{2},z_{3},z_{4}$, we have the complex version of Isserlis’s theorem:
(411) $\langle z_{1}z_{2}z_{3}z_{4}\rangle=\langle z_{1}z_{2}\rangle\langle z_{3}z_{4}\rangle+\langle z_{1}z_{3}\rangle\langle z_{2}z_{4}\rangle+\langle z_{1}z_{4}\rangle\langle z_{2}z_{3}\rangle,$
which follows from the real case using that the above is (with some abuse of terminology) a multilinear map (over $\mathbb{C}$). Explicitly, we can prove the case where $z_{m}$, $m\geq n$ are real by performing mathematical induction on $n$, where the induction step is done by separating real and imaginary parts of $z_{n}$.
@@ 14. Acknowledgments
Y.I.L. acknowledges support from the McGill University Department of Physics. The idea of ignoring quantitatively small deviations between experiment and theory is due to Paul François (Université de Montréal).